The demand and supply equations for a certain item are given by:
D = 5p + 40
S = -p^2 + 30p - 8
Find the equilibrium price..... How do I find the price when I am not sure of the actual numbers? Thank you!

Answer by stanbon(75887)

The demand and supply equations for a certain item are given by:
D = 5p + 40
S = -p^2 + 30p - 8
Find the equilibrium price.
------------
Equilibrium occurs when supply equal demand
EQUATION:
-p^2+30p-8=5p+40
-p^2+25p-48=0
p=[-25+-sqrt(625-4*-1*-48]/-2
p=[-25+-sqrt(625-192)]/-2
p=[-25+-sqrt(433)]/-2
p=2.09 or 22.90
-------
Cheers,
Stan H.